OpenStudy (anonymous):

The length of a rectangle is 2 feet more than twice its width. It's area is 84 square inches. Find the length and width of the rectangle

OpenStudy (confluxepic):

\[\huge \color{blue}{Welcome~To~OpenStudy}\]

OpenStudy (confluxepic):

A=lw

OpenStudy (confluxepic):

84=(2w+2)(w)

OpenStudy (confluxepic):

84=2w^2+2w

OpenStudy (confluxepic):

42=w^2+w

OpenStudy (confluxepic):

42-w=w^2